Denominator of fraction is 6 more then the numerator. If one is added to numerator and 3 is subtracted from the denominator it becomes 3/4. Find fraction.
Answer by Boreal(15235)

x/x+6
x+1/x+3, which is adding one to the numerator and subtracting 3 from the denominator,
it equals 3/4
cross-multiply
4(x+1)=3(x+3)
4x+4=3x+9
x=-5
5/11 becomes 6/8, which is 3/4
Fraction is 5/11